What is quantum tunneling? Calculate (i) the tunneling length and (ii) the tunneling probability when an electron of energy 4.0 eV is incident on a potential barrier of height 10 eV and width of 0.2 nm.
1
Expert's answer
2021-11-01T19:25:13-0400

Quantum tunnelling is the quantum mechanical phenomenon where a particle (more accurately, its wavefunction) can propagate through a potential barrier.


Let tunneling lenght be "a = 0.2nm = 0.2\times 10^{-9}m", the energy of the particle "E = 4eV \approx 6.409 \times 10^{-19}J" and the potential height "U = 10eV\approx 16.02\times 10^{-19}J". The tunneling propability is then given as follows (see https://courses.physics.illinois.edu/phys485/fa2015/web/tunneling.pdf, page 13):

"T = \left( 1 + \dfrac{\sinh^2(\beta a)}{4(E/V)(1-E/V)} \right)^{-1}"


where "\beta=\dfrac{\sqrt{2m(V-E)}}{\hbar}", "m = 9.1\times 10^{-31}kg" is the mass of electron and "\hbar=1.055\times 10^{-34}Js". Thus, obtain:


"\beta=\dfrac{\sqrt{2\cdot 9.1\times 10^{-31}kg\cdot (10-4)eV}}{1.055\times 10^{-34}Js} \approx12.54\times 10^{9}m^{-1}"


"T = \left( 1 + \dfrac{\sinh^2(12.54\times 10^{9}m^{-1}\cdot 0.2\times 10^{-9}m)}{4(4eV/10eV)(1-4eV/10eV)} \right)^{-1} \approx 0.025"

Answer. (i) 0.2nm, (ii) 0.025.